OpenStudy (cwrw238):

Positive integers x and y satisfy the equation sqrt x - sqrt 11 = sqrt y What is the maximum possible value of x/y? A 2 B 4 C 8 D 11 E 44 i spent a couple of hours on this and i think the answer is 4 but i'm far from sure.

OpenStudy (anonymous):

the integer part is throwing me off, but if you have only 4 choices should be easy enough to check which works right?

OpenStudy (anonymous):

oh never mind that is the output, not input sorry

OpenStudy (cwrw238):

yes - i did it by guessing and working out - not in a very formal way and figured that x = 44 and y = 11 giving x/y = 4

OpenStudy (anonymous):

on the other hand both x and y are integers right? that means you have very few choices for x. \[y=x-2\sqrt{11x}+11\] so your only choices for x are \[11,11^3,...\]

OpenStudy (cwrw238):

its deceptively difficult - for me anyway!

OpenStudy (anonymous):

oh right, what happens if you ignore the integer part?

OpenStudy (cwrw238):

sorry - dont know what you mean

OpenStudy (anonymous):

Nothing in the problem says x and y are integers.

OpenStudy (anonymous):

Stumped!!

OpenStudy (anonymous):

looks like 4 to me

OpenStudy (cwrw238):

x and y are positive integers

OpenStudy (cwrw238):

yea - i think it is - other options gave irrational values for x and y

OpenStudy (cwrw238):

if my arithmetic is correct - its worn me out!

OpenStudy (anonymous):

4 can't be right. Take x=16. y=0.467. Ratio is 34.26.

OpenStudy (anonymous):

yes i think 4 is it. if you try \[x=99\] you get \[y=44\] so \[\frac{x}{y}=\frac{9}{4}\]

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

@GT \[x=n^2\times 11, n \in \mathbb Z^+\] or else y is not an integer

OpenStudy (anonymous):

Sorry - i missed the first line of the problem. duh! :)

OpenStudy (cwrw238):

ahhh - right satellite

OpenStudy (cwrw238):

Thats ok GT

OpenStudy (anonymous):

ok that was a mistake but i think it is the right approach

OpenStudy (cwrw238):

how did you get - oh ok

OpenStudy (anonymous):

n^2 and (n-1)^2 are the values for coefficient of 11 for x and y. So, the value of x/y has to be n/(n-1). Maximum value is 2?

OpenStudy (anonymous):

\[x=n^2\times 11\] \[\sqrt{n^2\times 11}-\sqrt{11}=\sqrt{y}\] \[n\sqrt{11}-\sqrt{11}=\sqrt{y}\] \[(n-1)\sqrt{11}=\sqrt{y}\] \[11(n-1)^2=y\] \[\frac{x}{y}=\frac{n^2\times 11}{11(n-1)^2}\] \[\frac{x}{y}=\frac{n^2}{(n-1)^2}\]

OpenStudy (anonymous):

now your choices for n are 1 which gets you zero for y so that is undefined. n = 2 which gives you 4, and this function is decreasing for all n > 1, so that is your best bet. namely n = 2 and your max is 4

OpenStudy (cwrw238):

right - i see - but - excuse me if this is a silly question how did you get the first line? x = n^2 x 11?

OpenStudy (anonymous):

oh you gave it to me!

OpenStudy (anonymous):

you said x could be 11,44 , etc. so we know it has to be n^2*11 else the square root is not an integer

OpenStudy (cwrw238):

right - thanks a lot

OpenStudy (anonymous):

yw

OpenStudy (cwrw238):

this was a question on a UK maths challenge for ages 16-17. some question